# RayGauss: Volumetric Gaussian-Based Ray Casting for Photorealistic Novel View Synthesis

### v1

@@ 6.3 Ablation Studies
(Ours) Spherical Harmonics 33.69 0.972
@@
(Ours) Anisotropic Gaussian 34.34 0.974
@@
(Ours) Anisotropic Gaussian 34.34 0.974
(Ours) RGB 30.38 0.961
(Ours) Spherical Harmonics 33.69 0.972
(Ours) Spherical Gaussians 34.11 0.972
(Ours) Spherical Gaussians + Spherical Harmonics 34.34 0.974
@@
3D Gaussian Splatting: We first compare our method with Gaussian basis function and colorimetric parameters similar to 3D Gaussian Splatting (Spherical Harmonics) and observe that we achieve better results with an advantage of +0.3 dB, demonstrating that our ray-casting algorithm enables superior rendering quality than splatting while also avoiding rendering artifacts such as flickering. This can be observed in the supplementary videos, which show the flickering in Gaussian Splatting that is absent in our approach.
Basis function: We then conduct tests to confirm the choice of the Gaussian function. It can be observed that the anisotropic Gaussian yields better results than its isotropic counterpart, with an average difference of +0.9 dB. Various classical basis functions were also tested, including those with locally native support (such as the Bump function) or globally supported (such as the Inverse Quadratic function). Several showed promising results, notably the Wendland function. However, the anisotropic Gaussian performs best with at least a +0.25 dB difference (see Tab. 4). This justifies the choice of the anisotropic Gaussian as the basis function.
Colorimetric parameters: Then, we study the influence of the parameterization of the emitted radiance $c$, particularly with RGB parameters without directional dependence, 16 spherical harmonics (SH), 16 spherical Gaussians (SG), and the combination of 16 SH and SG functions yielding the best results: 9 SH (order 2) and 7 SG. We observe that parameterizations with directional dependence (SH, SG, and SH+SG) significantly improve quality. Furthermore, the parameterization providing the best quantitative results is the combination of SH and SG. This can be justified because SHs are robust for representing low-frequency phenomena, while SGs can adapt to high frequencies.
@@ 4 Optimization details
This section provides further details on the optimization parameters used in our experiments. We use the Adam gradient descent optimization algorithm [12]. We recall that the parameters optimized by our approaches are as follows: $P=\{(\tilde{\sigma}_{l},\tilde{c}_{l},\mu_{l},\mathbf{q}_{l},\mathbf{s}_{l})\mid l=1,\ldots,N\}$ where $\tilde{\sigma}_{l}$ is the density parameter of the $l$-th primitive, $\tilde{c}_{l}$ summarizes the colorimetric parameters: lobe sharpness $\lambda_{l,j}$ , lobe direction $\mathbf{p}_{l,j}$, coefficients $k_{l,j}$, for the $j-th$ Spherical Gaussian and coefficients $\tilde{c}_{l,jm}$ for Spherical Harmonic of degree $j$ and order $m$, while $\mu_{l}$, $\mathbf{q}_{l}$, and $\mathbf{s}_{l}$ are the mean position, quaternion, and scale parameter used to evaluate the basis function of the $l$-th primitive. The initial parameters are obtained from a point cloud created by Structure-From-Motion methods such as Colmap[27].
@@
We will now describe the main parameters of our optimization, starting with the learning rates. In the context of evaluations on the Blender dataset, we set the learning rates to $1.5\times 10^{-1}$ for density, $1.3\times 10^{-3}$ for constant color parameters (RGB), $1.1\times 10^{-4}$ for Spherical Harmonics coefficients, $6.0\times 10^{-4}$ for Spherical Gaussians coefficients, $1.0\times 10^{-1}$ for lobe sharpness, $2.0\times 10^{-3}$ for lobe direction, $1.2\times 10^{-2}$ for scale parameters, $3.0\times 10^{-4}$ for quaternions, and a learning rate with exponential decay starting from $1.5\times 10^{-5}$ and reaching $2.5\times 10^{-6}$ after 30,000 iterations for mean positions. We optimize the scene using a white background to place ourselves in evaluation conditions like other state-of-the-art methods. In the case of the Mip-NeRF360 dataset, we use similar learning rates except for the density, for which we use a learning rate with exponential decay starting from 0.5 and reaching 0.01 in 30,000 iterations. This allows us to accelerate the beginning of the optimization, particularly for outdoor scenes with a large number of Gaussians. Furthermore, we train with a black background on this dataset to place ourselves in evaluation conditions similar to Gaussian Splatting and current state-of-the-art methods [11][32].
@@
More generally, we optimize the scene by backpropagation, with the optimization taking 30,000 iterations (1 image per iteration). The process of Adaptive Gaussian Control is similar to that used in 3D Gaussian Splatting [11], periodically increasing, every $i_{densify}$ iterations, the number of Gaussians in the scene using a heuristic based on the condition $\nabla L_{\mu_{l}}>\nabla L_{\epsilon}$, where $\nabla L_{\mu_{l}}$ denotes the gradient on the mean position of the basis functions and $\nabla L_{\epsilon}$ the threshold from which densification is applied. In our case, we densify every 500 iterations starting from the 500th iteration up to the 15,000th iteration, and we set $\nabla L_{\epsilon}=0.00002$ for the Mip-NeRF 360 dataset and $\nabla L_{\epsilon}=0.00005$ for the Blender dataset. Furthermore, we remove primitives whose density parameter $\tilde{\sigma}$ is below a threshold, set to 0.1 for our experiments on the Blender dataset and 0.01 on the Mip-NeRF 360 dataset. Finally, as described in the main paper, we gradually unlock the colorimetric parameters from the lowest frequency representation to the potentially higher frequency one: harmonics of degree 0, 1, 2, and finally, 7 spherical Gaussians, resulting in 16 functions to represent the color of a primitive. The unlocking of these parameters occurs every 1000 iterations. Also, one can refer to Algorithm 3 to gain a broader view of the optimization process.
@@
Furthermore, it should be noted that the number of parameters in our representation is $87N_{P}$, where $N_{P}$ is the number of primitives. Specifically, we have the following parameter counts: 3 for the mean position $\mu$, 3 for the scale parameter $\mathbf{s}$, 4 for the quaternions $\mathbf{q}$, 1 for the density parameter $\tilde{\sigma}$, 27 for spherical harmonic coefficients, 21 for spherical Gaussian coefficients, 7 for lobe sharpness parameters, and 21 for lobe direction parameters. Moreover, as mentioned earlier, our current implementation requires allocating a buffer of size $N_{max,P}\times N_{ray}$, where $N_{max,P}$ is the maximum number of primitives per ray per slab allowed, and $N_{ray}$ the number of rays launched. Therefore, the limiting factor of our implementation in terms of memory mainly depends on the allocation of these data.
@@ 6 OptiX API applications
7 Study of two main parameters of the method
@@
0.000125 34.34 46 min 17.4 FPS
0.00025 34.34 32 min 25.8 FPS
0.0005 34.23 24 min 42.8 FPS
0.001 33.56 21 min 54.7 FPS
Table 5: Study of the influence of the parameter $\Delta t$ on the Blender dataset. $\Delta t$ is the distance between two samples along a ray. Values are average per scene. In gray, the parameter used in other experiments. Rendering time is for 800x800 pixels image.
Tab. 6 studies the influence of the parameter $\sigma_{\epsilon}$ on the Blender dataset. The results come from training on the Blender Dataset with PSNR averaged over the 8 scenes of this synthetic dataset. We can see that increasing $\sigma_{\epsilon}$ speeds up the training and rendering times of the method. On the contrary, by decreasing $\sigma_{\epsilon}$, we increase the size of the Gaussians when calculating the intersections with the rays, which increases the training and rendering times, while improving the quality of rendering.
@@ 9 Detailed results
Tab. 9 and Tab. 10 show the detailed results of the main paper with metrics PSNR, SSIM, and LPIPS on the Blender and Mip-NeRF 360 datasets. Some methods have no available code, so we were not able to report information about SSIM and LPIPS (for example, for PointNet++).
@@
RayGauss (ours) 0.010 0.031 0.011 0.016 0.013 0.027 0.004 0.092 0.026
Table 9: PSNR, SSIM and LPIPS scores on the Blender dataset. All methods are trained on the train set with full-resolution images (800x800 pixels) and evaluated on the test set with full-resolution images (800x800 pixels).
@@
Table 10: PSNR, SSIM and LPIPS scores on the Mip-NeRF 360 dataset. All methods are trained and tested on downsampled images by a factor of 8.

### v2

@@ 6.3 Ablation Studies
(Ours) Spherical Harmonics 33.75 0.972
@@
(Ours) Anisotropic Gaussian 34.53 0.974
@@
(Ours) Anisotropic Gaussian 34.53 0.974
(Ours) RGB 30.39 0.961
(Ours) Spherical Harmonics 33.75 0.972
(Ours) Spherical Gaussians 34.40 0.973
(Ours) Spherical Gaussians + Spherical Harmonics 34.53 0.974
@@
3D Gaussian Splatting: We first compare our method with Gaussian basis function and colorimetric parameters similar to 3D Gaussian Splatting (Spherical Harmonics) and find that we achieve a +0.36 dB improvement, demonstrating that our ray-casting algorithm offers superior rendering quality while avoiding splatting artifacts, as seen in the supplementary videos, which highlights the absence of flickering and splatting artifacts compared to Gaussian Splatting.
Basis function: We then conduct tests to confirm the choice of the Gaussian function. Results show the anisotropic Gaussian outperforms the isotropic one by +1.09 dB on average. Various classical basis functions were also tested, including those with locally native support (such as the Bump function) or globally supported (such as the Inverse Quadratic function). Several showed promising results, notably the Wendland function. However, the anisotropic Gaussian performs best with at least a +0.35 dB difference (see Tab. 4). This justifies the choice of the anisotropic Gaussian as the basis function.
Colorimetric parameters: We analyze the impact of different radiance $c$ parameterizations: RGB without directional dependence, 16 spherical harmonics (SH), 16 spherical Gaussians (SG), and the combination of 16 SH and SG functions yielding the best results: 9 SH (order 2) and 7 SG. We observe that parameterizations with directional dependence (SH, SG, and SH+SG) significantly improve quality. Then, the parameterization providing the best quantitative results is the combination of SH and SG. This can be justified because SHs are robust for representing low-frequency phenomena, while SGs can adapt to high frequencies.
@@ 4 Optimization details
This section provides further details on the optimization parameters used in our experiments. We use the Adam gradient descent optimization algorithm [13]. We recall that the parameters optimized by our approaches are as follows: $P=\{(\tilde{\sigma}_{l},\tilde{c}_{l},\mu_{l},\mathbf{q}_{l},\mathbf{s}_{l})\mid l=1,\ldots,N\}$ where $\tilde{\sigma}_{l}$ is the density parameter of the $l$-th primitive, $\tilde{c}_{l}$ summarizes the colorimetric parameters: lobe $\lambda_{l,j}$ , lobe direction $\mathbf{p}_{l,j}$, coefficients $k_{l,j}$, for the $j-th$ Spherical Gaussian and coefficients $\tilde{c}_{l,jm}$ for Spherical Harmonic of degree $j$ and order $m$, while $\mu_{l}$, $\mathbf{q}_{l}$, and $\mathbf{s}_{l}$ are the mean position, quaternion, and scale parameter used to evaluate the basis function of the $l$-th primitive. The initial parameters are obtained from a point cloud created by Structure-From-Motion methods such as Colmap[28].
@@
We will now describe the main parameters of our optimization, starting with the learning rates. In the context of evaluations on the Blender dataset, we employed a learning rate for density with exponential decay, starting at $5\times 10^{-1}$ and decreasing to $3\times 10^{-2}$ over 30,000 iterations. The learning rates for other parameters were set as follows: $1.0\times 10^{-3}$ for constant color parameters (RGB), $2.6\times 10^{-4}$ for Spherical Harmonics coefficients, $3.6\times 10^{-4}$ for Spherical Gaussians coefficients, $4.5\times 10^{-2}$ for lobe sharpness, $3.0\times 10^{-3}$ for lobe direction, $1.2\times 10^{-2}$ for scale parameters, $2.2\times 10^{-4}$ for quaternions, and a learning rate with exponential decay starting from $1.7\times 10^{-5}$ and reaching $1.0\times 10^{-6}$ after 30,000 iterations for mean positions. We optimize the scene using a white background to place ourselves in evaluation conditions like other state-of-the-art methods. In the case of the Mip-NeRF360 dataset, we use similar learning rates except for the density, for which we use a learning rate with exponential decay starting from $5\times 10^{-1}$ and reaching $1\times 10^{-4}$ in 30,000 iterations. This allows us to accelerate the beginning of the optimization, particularly for outdoor scenes with a large number of Gaussians. Furthermore, we train with a black background on this dataset to place ourselves in evaluation conditions similar to Gaussian Splatting and current state-of-the-art methods [12][33].
@@
More generally, we optimize the scene by backpropagation, with the optimization taking 30,000 iterations (1 image per iteration). Quaternions are renormalized after each optimization step to maintain a valid rotation matrix representation. The process of Adaptive Gaussian Control is similar to that used in 3D Gaussian Splatting [12], periodically increasing, every $i_{densify}$ iterations, the number of Gaussians in the scene using a heuristic based on the condition $\nabla L_{\mu_{l}}>\nabla L_{\epsilon}$, where $\nabla L_{\mu_{l}}$ denotes the gradient on the mean position of the basis functions and $\nabla L_{\epsilon}$ the threshold from which densification is applied. In our case, we densify every 500 iterations for Mip-NeRF360 and 300 iterations for Blender dataset, starting from the 500th iteration up to the 15,000th iteration, and we set $\nabla L_{\epsilon}=0.00002$ for the Mip-NeRF 360 dataset and $\nabla L_{\epsilon}=0.00004$ for the Blender dataset. Furthermore, we remove primitives whose density parameter $\tilde{\sigma}$ is below a threshold, set to 0.1 for our experiments on the Blender dataset and 0.01 on the Mip-NeRF 360 dataset. Finally, as described in the main paper, we gradually unlock the colorimetric parameters from the lowest frequency representation to the potentially higher frequency one: harmonics of degree 0, 1, 2, and finally, 7 spherical Gaussians, resulting in 16 functions to represent the color of a primitive. The unlocking of these parameters occurs every 1000 iterations. We can specify here that the spherical harmonic and spherical gaussian coefficients, as well as the lobe sharpness, are initialized to zero, while the lobe axes are randomly initialized on the unit sphere and constrained to it during optimization. Also, one can refer to Algorithm 3 to gain a broader view of the optimization process.
@@
Furthermore, it should be noted that the number of parameters in our representation is $87N_{P}$, where $N_{P}$ is the number of primitives. Specifically, we have the following parameter counts: 3 for the mean position $\mu$, 3 for the scale parameter $\mathbf{s}$, 4 for the quaternions $\mathbf{q}$, 1 for the density parameter $\tilde{\sigma}$, 27 for spherical harmonic coefficients, 21 for spherical Gaussian coefficients, 7 for lobe parameters, and 21 for lobe direction parameters. Moreover, as mentioned earlier, our current implementation requires allocating a buffer of size $N_{max,P}\times N_{ray}$, where $N_{max,P}$ is the maximum number of primitives per ray per slab allowed, and $N_{ray}$ the number of rays launched. Therefore, the limiting factor of our implementation in terms of memory mainly depends on the allocation of these data.
@@ 6 OptiX API applications
7 Analysis of $\Delta t$ and $\sigma_{\epsilon}$
@@
0.00125 34.53 46 min 17.4 FPS
0.0025 34.53 32 min 25.8 FPS
0.005 34.52 24 min 42.8 FPS
0.01 33.90 20 min 50.1 FPS
Table 5: Study of the influence of the parameter $\Delta t$ on the Blender dataset. $\Delta t$ is the distance between two samples along a ray. Values are averaged per scene. In gray, the parameter used in other experiments. Rendering time is for 800x800 pixels image.
Tab. 6 studies the influence of the parameter $\sigma_{\epsilon}$ on the Blender dataset. The results come from training on the Blender Dataset with PSNR averaged over the 8 scenes of this synthetic dataset. We can see that increasing $\sigma_{\epsilon}$ speeds up the training and rendering times of the method. On the contrary, by decreasing $\sigma_{\epsilon}$, we increase the size of the Gaussians when calculating the intersections with the rays, which increases the training and rendering times, while improving the quality of rendering. The gray line also corresponds to the choice of the parameter $\sigma_{\epsilon}$ for all other experiments on Blender dataset (main article and supplementary), for Mip-NeRF360, we use $\sigma_{\epsilon}=0.01$.
@@ 9 Detailed results
Tab. 9 and Tab. 10 show the detailed results of the main paper with metrics PSNR, SSIM, and LPIPS on the Blender and Mip-NeRF 360 datasets. Some methods have no available code, so we were not able to report information about SSIM and LPIPS (for example, for PointNeRF++).
@@
NeuRBF[6]* 0.016 0.061 0.016 0.021 0.015 0.032 0.008 0.114 0.035
RayGauss (ours) 0.009 0.030 0.011 0.015 0.012 0.026 0.004 0.088 0.024
Table 9: PSNR, SSIM and LPIPS (with VGG network) scores on the Blender dataset. All methods are trained on the train set with full-resolution images (800x800 pixels) and evaluated on the test set with full-resolution images (800x800 pixels).
@@
Table 10: PSNR, SSIM and LPIPS (with VGG network) scores on the Mip-NeRF 360 dataset. All methods are trained and tested on downsampled images by a factor of 8.